The Voyager I spacecraft left Earth in 1977 traveling at about 17 km/s. If it was headed in the right direction, in what year would it reach the nearest star, Proxima Centauri at a distance of 1.3 pc?
Answer #1

We know that

time = distance/Velocity

Velocity = 17 km/sec = 17000 m/sec

distance between proxima centauri and earth = 1.3 pc

Since 1 pc = 1 parsec = 3.086*10^16 m

So,

time = 1.3*3.086*10^16 m/(17000 m/sec)

time = 2.36*10^12 sec

Now we know that 1 yr = 3.154*10^7 sec

So,

time = 2.36*10^12 sec*(1 yr/(3.154*10^7 sec)) = 74785 yr

Since currently we are in 1977, So the year it will reach Proxima centauri will be = 1977 + 74785 = 76762 yr

Please Upvote.